# REGen: A Reliable Evaluation Framework for Generative Event Argument Extraction

### v1

@@ Abstract
Event argument extraction identifies arguments for predefined event roles in text. Traditional evaluations rely on exact match (EM), requiring predicted arguments to match annotated spans exactly. However, this approach fails for generative models like large language models (LLMs), which produce diverse yet semantically accurate responses. EM underestimates performance by disregarding valid variations, implicit arguments (unstated but inferable), and scattered arguments (distributed across a document). To bridge this gap, we introduce Reliable Evaluation framework for Generative event argument extraction (REGen), a framework that better aligns with human judgment. Across six datasets, REGen improves performance by an average of 23.93 F1 points over EM. Human validation further confirms REGen’s effectiveness, achieving 87.67% alignment with human assessments of argument correctness.
@@ 1 Introduction
Previous works have attempted to address these issues using embedding-based relaxed matching, which considers two arguments similar if they have high embedding similarity Han et al. (2024). However, this approach fails to capture semantically similar arguments with different lexical forms and wrongly classifies arguments with high token overlap as similar Sharif et al. (2024). For example, in Figure 2 for the role ‘patient concerns’, the ground-truth argument ‘limited insurance coverage’ and the predicted argument ‘coverage limitations for FLA and cryotherapy’ refers to the same issue. Due to lexical variation, relaxed matching fails to capture this. In contrast, consider a role ‘date’ for which ground-truth and predicted arguments are ‘18 April 2024’ and ‘20 April 2024’, respectively. These two arguments are different, but relaxed matching considers them the same due to high token overlap. Context is needed when evaluating these arguments. Recent work by Lu et al. (2024) used LLMs as judges to identify similar arguments. This approach requires a large number of inferences, adding significant computational costs. Additionally, without human validation, LLM-based judgments can produce unreliable results. Relying solely on relaxed match or judge-based approaches can overestimate performance by incorrectly classifying non-matching arguments as matches, leading to inflated and unreliable model assessments.
@@
After filtering out exact and relaxed matches, unmatched arguments are carried forward for complex matching. The CM stage captures semantically similar arguments based on context despite lexical and/or syntactic differences. We leverage LLM as a judge Zheng et al. (2023) at this stage for determining argument similarity. Finally, in the judgment alignment stage, we propose a novel Judgment Aligned Match (JAM) score to factor in the scores from each level to account for misjudgments based on extensive human validation. This framework ensures evaluation accuracy, cost-effectiveness, and better alignment with human judgments.
@@
Reproducibility: Our code, evaluation framework, the judge and alignment datasets, and other relevant resources are available at https://github.com/omar-sharif03/EAE-Eval.
@@ 2.2 Level-1: Exact Match
An exact match (EM) pair is defined when $p_{i}=g_{j}$, forming a list of EM pairs such as $[(p_{1},g_{2}),(p_{3},g_{5}),\ldots,(p_{x},g_{y})]$. Precision and recall for the EM level are computed as:
@@ 2.4 Level 3: Complex Match
[…]
@@ 2.5 Alignment with Human Judgments
We first calculate the deviation rate of a matching model $(M)$ on a dataset $(DT)$ by measuring the number of disagreements between the model and the human evaluator. The deviation rate is computed using equation 3.
@@ 3.2 Judge Selection for Complex Match
Judge data annotation: We construct a judge dataset comprising 900 argument pairs to select the best judge model. Specifically, we randomly select 150 pairs not matched under exact or relaxed criteria from each of the six datasets evaluated. Each pair is annotated as ‘match’ or ‘non-match’ by a human annotator. A second human verifies the labels, and disagreements are resolved through discussion to finalize the annotations.
Judge LLM selection: We evaluate both open-source (Llama3.1-70B) and closed-source (GPT-4o, GPT-4o-mini, GPT-3.5) models as potential judges, assessing their performance in zero-shot and chain-of-thought settings. GPT-4o, with a zero-shot prompt, achieves the highest agreement with human judgments, scoring 86.17. Therefore, we selected GPT-4o as the judge model for complex match evaluation. Note that the choice of judge is orthogonal to our proposed framework. The selected judge model can easily be swapped with newer or better alternatives without further modifications. Appendix C provides additional details on judge selection and relevant prompts.
@@ 3.3 Judgment Alignment
Among the datasets, the PHEE dataset showed the highest alignment (92.67%) with human judgments, while DocEE had the lowest (80.67%). On average, the REGen framework achieved 87.67% alignment with human evaluators across all six datasets and matching levels. Our analysis reveals primary reasons for judgment disagreements are (1) Numerical nuances: the model often failed to distinguish numerical differences. Such as for a role ‘drug-dosage,’ it incorrectly treated ‘14 mg’ and ‘6 mg’ as equivalent. (2) Temporal variations: dates such as ‘18 April’ versus ‘20 April’ or days like ‘Thursday’ versus ‘Friday’ were incorrectly judged as similar. (3) Coreference handling: datasets like RAMS and WikiEvents frequently used pronouns (e.g., ‘he’, ‘they’) in the ground truth, while models predicted specific names (e.g., ‘John’). This mismatch led to judgment errors, especially when documents contained multiple names, confusing the model.
@@ 4.1 Datasets and Experimental Setup
We used six standard EAE datasets from diverse domains to evaluate REGen. These datasets include: RAMS Ebner et al. (2020) (news), GENEVA Parekh et al. (2023) (book, news, journal articles), DocEE Tong et al. (2022) (long news documents), WikiEvents Li et al. (2021) (Wikipedia texts), DiscourseEE Sharif et al. (2024) (online health discourse), and PHEE Sun et al. (2022) (pharmacovigilance texts). Prior works, such as Huang et al. (2024) and Lu et al. (2024) have evaluated LLMs using small test subsets sampled and merged from multiple datasets. Thus not reflecting actual performance of LLMs on these datasets. We conduct evaluations using the complete official test sets of the selected datasets to provide a more reliable assessment of LLMs’ performance on these benchmarks. Detailed statistics for these test datasets are presented in Table 3. Appendix A contains additional details on the data preparation steps.
@@ 4.2 EAE Models
LLM Based Models: We perform comprehensive experiments using open-source and closed-source LLMs from different model families of various parameter sizes, including Phi-3.5 (3.8B), Gemma-1.1 (7B), Mixtral (8x7B), Llama-3.1 (70B), and GPT-4o. We evaluate all the models in two prompt settings: zero-shot and chain-of-thought. We employed question-guided prompting as previous works achieved SOTA performance using this approach Lu et al. (2023); Hsu et al. (2022); Du and Cardie (2020). Specifically, models are prompted with (Instruction, Document, Question) to generate $\rightarrow$ (Arguments), where each question is tailored to extract specific role [^1] Role-specific questions for each dataset can be found here: https://tinyurl.com/48y2wus3. Sample questions for the datasets are presented in Table 13.
Different LLMs require prompts and in-context samples tailored to each model and dataset. In practice, users select the optimal prompt using a trial-and-error approach Ziems et al. (2024); Zamfirescu-Pereira et al. (2023). However, in our experiments, iterating over various prompts to find the optimal prompt for each model and dataset is impractical. Instead, we opted to use a consistent prompt across all models and datasets to (i) ensure a fair comparison among the models and (ii) eliminate the confounding factors related to prompt optimization. Generic templates for zero-shot and chain-of-thought prompts for argument extraction are illustrated in Figures 8 and 9, respectively. Additional descriptions of the models are provided in Appendix B.
@@ 5 Results
Table 2: Comparison of average F1-scores of the LLMs between Exact Match (EM) and REGen evaluation frameworks.
@@
Table 4: Evaluation results using the REGen framework for event argument extraction across the six datasets. The table reports F1-scores for models assessed at different evaluation levels: Exact Match (EM), Relaxed Match (RM), Complex Match (CM), and Judgment-Aligned Match (JAM). Due to space constraints, detailed precision, recall, and F1-scores are provided in Appendix Tables 7-12. The highest and the second-highest values in a column are highlighted using a dark shade and light shade, respectively.
Significant improvement in F1-score across all datasets: Table 4 illustrates performance of various models using REGen framework. We observed a notable performance boost when models transitioned from the EM to the JAM score. For instance, the F1-score for the top-performing GPT-4o model increased from 16.82 with EM to 46.16 with JAM in the DiscourseEE dataset. Additionally, the average F1-scores of the LLMs shown in Table 2 exhibit that all evaluated datasets achieved considerable performance gains, averaging 23.93 points. The increase in F1 score for the GENEVA dataset was 32.49, representing a 238.52% improvement over the standard EM evaluation. Similar substantial gains were noted in other datasets, such as 26.71 for DiscourseEE and 24.32 for DocEE.
@@
REGen framework is more reliable (87.67% alignment): REGen shows no/minimal errors in the performance under exact and relaxed match scoring. While there is some overestimation due to misjudgments in the complex match step, our extensive validation indicates an 87.67% alignment with human judgments (see Table 1). The JAM score incorporates this human alignment, ensuring the overall reliability of the framework. Additionally, the reported scores are more explainable, as they include a clear breakdown of performance gains at each level (EM, RM, CM, and JAM).
@@
Recall is on average higher than precision in all settings: Our fine-grained analysis (see Tables 7 to 12) reveals LLMs achieve higher recall than precision. Such as the GPT-4o model in the DocEE dataset achieved a JAM recall of 68.41 compared to a precision of only 42.12. This indicates while the models are effective in identifying ground-truth arguments, they tend to over-predict, impacting the overall F1-score. In this work, we used a single prompt for all the models and datasets, which might have contributed to this overprediction. Future research should focus on pushing the performance through dataset- and model-specific prompting to enhance precision without sacrificing recall.
@@ 6 Related Work
Generative Event Argument Extraction: Early studies on event argument extraction (EAE) treated it as an extractive or token-level classification task Doddington et al. (2004); Du and Cardie (2020). These efforts primarily focused on identifying argument spans directly found in the text Sun et al. (2022). Recently, EAE has been formulated as a generative task where pre-trained language models are guided with natural language to fill templates or generate arguments Hsu et al. (2022). Sharif et al. (2024) argue that this generative formulation better suits real-world applications as it can capture implicit and scattered arguments better. With the emergence of LLMs, generative model-based argument extraction gained more traction Sun et al. (2024); He et al. (2024). So, we focus on generative extraction covering diverse models and datasets.
Evaluations for Generative EAE: Existing works for generative EAE primarily rely on exact matching for evaluation Huang et al. (2024). This strict approach unfairly penalizes models, even when the generated output is correct. To address this, Han et al. (2024) adopt a relaxed matching approach, considering arguments similar if their embedding-based similarly exceeds a threshold of 0.5. Similarly, Sharif et al. (2024) used a threshold of 0.75. However, this approach has limitations. It fails to capture semantically similar arguments with different lexical or syntactic forms and wrongly classifies arguments with high token overlap as similar. Thus, performance reported solely on relaxed matching is unreliable. More recently, Lu et al. (2024) employed LLMs to determine argument similarity. Nonetheless, this approach incurs significant computational overhead and demands extensive human validation. Our REGen framework combines the strengths of exact, relaxed, and LLM-based matching. It systematically reduces misjudgments, computational costs, and the need for human validation.
@@ 7 Conclusion
This paper presents REGen, a novel evaluation framework for EAE. Our extensive experiments and human validation demonstrate its effectiveness, with a 23.93 points improvement in average F1 score across six EAE datasets and reliability, with 87.67% alignment with human judgments. We highlight the limitations of current evaluation approaches and illustrate how REGen addresses these issues. Furthermore, our analysis reveals that previous studies have underestimated the true performance of LLMs. We believe that REGen fills a critical gap in EAE research and motivates future work to explore the generative model’s capability in solving other information extraction tasks, e.g. relation extraction, entity extraction, and beyond.
@@ 8 Limitations
Intended Use:
We will release our judge and alignment datasets upon the acceptance of the paper to facilitate future research on generative argument extraction evaluation.
Annotation:
Judge and alignment data annotation were conducted by trained NLP researchers. All annotators were compensated as per the standard paying rate of the author’s institution. Key characteristics of our annotators include: (a) graduate students, (b) 3-6 years of research experience, and (c) a mix of native and non-native English speakers. We provided annotators detailed annotation guidelines, including argument extraction, semantical similarity, and the type of information we wanted to compare to mitigate potential biases.
Reproducibility
Details on models and dataset processing are provided in Appendices A and B. The evaluation framework, code, and processed datasets are available at https://github.com/omar-sharif03/EAE-Eval.

### v2

@@ Abstract
Event argument extraction identifies arguments for predefined event roles in text. Existing work evaluates this task with exact match (EM), where predicted arguments must align exactly with annotated spans. While suitable for span-based models, this approach falls short for large language models (LLMs), which often generate diverse yet semantically accurate arguments. EM severely underestimates performance by disregarding valid variations. Furthermore, EM evaluation fails to capture implicit arguments (unstated but inferable) and scattered arguments (distributed across a document). These limitations underscore the need for an evaluation framework that better captures models’ actual performance.
To bridge this gap, we introduce REGen, a Reliable Evaluation framework for Generative event argument extraction. REGen combines the strengths of exact, relaxed, and LLM-based matching to better align with human judgment. Experiments on six datasets show that REGen reveals an average performance gain of +23.93 F1 over EM, reflecting capabilities overlooked by prior evaluation. Human validation further confirms REGen’s effectiveness, achieving 87.67% alignment with human assessments of argument correctness.
@@ 1 Introduction
Previous works have attempted to address these issues using embedding-based relaxed matching, which considers two arguments similar if they have high embedding similarity Han et al. (2024). However, this approach fails to capture semantically similar arguments with different lexical forms and wrongly classifies arguments with high token overlap as similar Sharif et al. (2024). For example, in Figure 2 for the role ‘patient concerns’, the ground-truth argument ‘limited insurance coverage’ and the predicted argument ‘coverage limitations for FLA and cryotherapy’ refer to the same issue. Due to lexical variation, relaxed matching fails to capture this. In contrast, consider a role ‘date’ for which ground-truth and predicted arguments are ‘18 April 2024’ and ‘20 April 2024’, respectively. These two arguments are different, but relaxed matching considers them the same due to high token overlap. Context is needed when evaluating these arguments. Recent work by Lu et al. (2024) used LLMs as judges to identify similar arguments. This approach requires a large number of inferences, adding significant computational costs. Additionally, without human validation, LLM-based judgments can produce unreliable results. Relying solely on relaxed match or judge-based approaches can overestimate performance by incorrectly classifying non-match arguments as matches, leading to inflated and unreliable model assessments. A detailed analysis of argument correctness by each method is shown in Table 6.
@@
After filtering out exact and relaxed matches, unmatched arguments are carried forward for complex matching. The CM stage captures semantically similar arguments based on context despite lexical and/or syntactic differences. We leverage LLM as a judge Zheng et al. (2023) to determine argument similarity. Finally, in the judgment alignment stage, we propose a novel Judgment Aligned Match (JAM) score to factor in the scores from each level to account for misjudgments based on human validation. This framework ensures evaluation accuracy, cost-effectiveness, and better alignment with human judgments.
@@
Reproducibility: Our code, evaluation framework, the judge and alignment datasets, and other relevant resources are available at https://github.com/Omar-Sharif/REGen.
@@ 2.2 Level-1: Exact Match
An exact match (EM) pair is defined when $p_{i}=g_{j}$, forming a list of EM pairs such as $(p_{1},g_{2}),(p_{3},g_{5}),\ldots,(p_{x},g_{y})]$. Precision and recall for the EM level are computed as:
@@ 2.4 Level 3: Complex Match
$\text{F1}_{l}=\frac{2*\text{P}_{l}*\text{R}_{l}}{\text{P}_{l}+\text{R}_{l}}$ (2)
@@ 2.5 Alignment with Human Judgments
We first calculate the deviation rate of a matching model $(M)$ on a dataset $(DT)$ by measuring the number of disagreements between the model and the human evaluator. The deviation rate is computed using equation 3. Addition details on the deviation rate or the alignment calculation are provided in Appendix B.
@@ 3.2 Judge Selection for Complex Match
Judge data annotation: We construct a judge dataset comprising 900 argument pairs (150 pairs per dataset) to select the best judge model. Specifically, we randomly select pairs not matched under exact or relaxed criteria, meaning they inherently represent challenging or ambiguous cases. Each pair is annotated as ‘match’ or ‘non-match’ by a human annotator. A second human verifies the labels, and disagreements are resolved through discussion to finalize the annotations.
Judge LLM selection: We evaluate both open-source (Llama3.1-70B) and closed-source (GPT-4o, GPT-4o-mini, GPT-3.5) models as potential judges, assessing their performance in zero-shot and chain-of-thought settings. GPT-4o, with a zero-shot prompt, achieves the highest agreement with human judgments, scoring 86.17. Therefore, we selected GPT-4o as the judge model for complex match evaluation. Note that the choice of judge is orthogonal to our proposed framework. The selected judge model can easily be swapped with newer or better alternatives without further modifications. Appendix E provides additional details on judge selection and relevant prompts.
@@ 3.3 Judgment Alignment
Among the datasets, the PHEE dataset showed the highest alignment (92.67%) with human judgments, while DocEE had the lowest (80.67%). On average, the REGen framework achieved 87.67% alignment with human evaluators. Our analysis reveals primary reasons for judgment disagreements are (1) Numerical nuances: the model often failed to distinguish numerical differences. Such as for a role ‘drug-dosage,’ it incorrectly treated ‘14 mg’ and ‘6 mg’ as equivalent. (2) Temporal variations: dates such as ‘18 April’ versus ‘20 April’ or days like ‘Thursday’ versus ‘Friday’ were incorrectly judged as similar. (3) Coreference handling: datasets like RAMS and WikiEvents frequently used pronouns (e.g., ‘he’, ‘they’) in the ground truth, while models predicted specific names (e.g., ‘John’). This mismatch led to judgment errors, especially when documents contained multiple names, confusing the model. We identified a total of 111 disagreements. Of these, 15 cases were due to numerical nuances, 10 cases involved temporal variations, 57 cases were related to coreference handling, and 29 cases were due to other issues, such as the model incorrectly matching unrelated arguments. A detailed breakdown of these disagreement categories for each dataset is provided in Table 2. Additional error analysis is provided in Appendix A.
Disagreement category
Datasets Numerical Temporal Coreference Other
DiscourseEE 7 1 15 1
PHEE 1 1 4 5
RAMS 0 0 10 5
GENEVA 0 0 8 7
DocEE 6 8 12 3
WikiEvents 1 0 8 8
Total 15 10 57 29
Table 2: Detailed breakdown of disagreement cases between human and judge model in evaluated EAE datasets.
@@ 4.1 Datasets and Experimental Setup
We used six standard EAE datasets from diverse domains to evaluate REGen. These datasets include: RAMS Ebner et al. (2020) (news), GENEVA Parekh et al. (2023) (book, news, journal articles), DocEE Tong et al. (2022) (long news documents), WikiEvents Li et al. (2021) (Wikipedia texts), DiscourseEE Sharif et al. (2024) (online health discourse), and PHEE Sun et al. (2022) (pharmacovigilance texts). Prior works, such as Huang et al. (2024) and Lu et al. (2024) have evaluated LLMs using small test subsets sampled and merged from multiple datasets. Thus not reflecting actual performance of LLMs on these datasets. We conduct evaluations using the complete official test sets of the selected datasets to provide a more reliable assessment of LLMs’ performance on these benchmarks. Detailed statistics for these test datasets are presented in Table 4. Appendix C contains additional details on the data preparation steps.
@@ 4.2 EAE Models
LLM Based Models: We perform comprehensive experiments using open-source and closed-source LLMs from different model families of various parameter sizes, including Phi-3.5 (3.8B), Gemma-1.1 (7B), Mixtral (8x7B), Llama-3.1 (70B), and GPT-4o. We evaluate all the models in two prompt settings: zero-shot and chain-of-thought. We employed question-guided prompting as previous works achieved SOTA performance using this approach Lu et al. (2023); Hsu et al. (2022); Du and Cardie (2020). Specifically, models are prompted with (Instruction, Document, Question) to generate $\rightarrow$ (Arguments), where each question is tailored to extract specific role [^1] Role-specific questions for each dataset can be found here: https://tinyurl.com/38en8e94. Sample questions for the datasets are presented in Table 18.
Different LLMs require prompts and in-context samples tailored to each model and dataset. In practice, users select the optimal prompt using a trial-and-error approach Ziems et al. (2024); Zamfirescu-Pereira et al. (2023). However, in our experiments, iterating over various prompts to find the optimal prompt for each model and dataset is impractical. Instead, we opted to use a consistent prompt across all models and datasets to (i) ensure a fair comparison among the models and (ii) eliminate the confounding factors related to prompt optimization. Generic templates for zero-shot and chain-of-thought prompts for argument extraction are illustrated in Figures 8 and 9, respectively. Additional descriptions of the models are provided in Appendix D.
@@ 5 Results
Table 3: Average F1-scores of LLMs in zero-shot and chain-of-thought settings, comparing Exact Match (EM) and REGen evaluation frameworks. Additional comparisons with Relaxed Match (RM) and Complex Match (CM) are reported in Table 9.
@@
Table 5: Evaluation results using the REGen framework for event argument extraction across the six datasets. The table reports F1-scores for models assessed at different evaluation levels: Exact Match (EM), Relaxed Match (RM), Complex Match (CM), and Judgment-Aligned Match (JAM). Due to space constraints, detailed precision, recall, F1-scores, and additional results are provided in Appendix Tables 10-17. The highest and the second-highest values in a column are highlighted using a dark shade and light shade, respectively.
Significant improvement in F1-score across all datasets: Table 5 illustrates performance of various models using REGen framework. We observed a notable performance boost when models transitioned from the EM to the JAM score. For instance, the F1-score for the top-performing GPT-4o model increased from 16.82 with EM to 46.16 with JAM in the DiscourseEE dataset. Additionally, the average F1-scores of all the LLMs shown in Table 3 exhibit that all evaluated datasets achieved considerable performance gains, averaging 23.93 points. The increase in F1 score for the GENEVA dataset was 32.49, representing a 238.52% improvement over the standard EM evaluation. Similar substantial gains were noted in other datasets, such as 26.71 for DiscourseEE and 24.32 for DocEE.
@@
REGen framework is more reliable (87.67% alignment): REGen shows no/minimal errors in the performance under exact match and relaxed match scoring. While there is some overestimation due to misjudgments in the complex match step, our extensive validation indicates an 87.67% alignment with human judgments (see Table 1). The JAM score incorporates this human alignment, ensuring the overall reliability of the framework. Additionally, the reported scores are more explainable, as they include a clear breakdown of performance gains at each matching level (EM, RM, CM, and JAM).
@@
Recall is on average higher than precision in all settings: Our fine-grained analysis (see Tables 12 to 17) reveals LLMs achieve higher recall than precision. Such as the GPT-4o model in the DocEE dataset achieved a JAM recall of 68.41 compared to a precision of only 42.12. This indicates while the models are effective in identifying ground-truth arguments, they tend to over-predict, impacting the overall F1-score. In this work, we used a single prompt for all the models and datasets, which might have contributed to this overprediction. Future research should focus on pushing the performance through dataset- and model-specific prompting to enhance precision without sacrificing recall.
@@ 6 Related Work
Generative Event Argument Extraction: Early studies on event argument extraction (EAE) treated it as an extractive or token-level classification task Doddington et al. (2004); Du and Cardie (2020). These efforts primarily focused on identifying argument spans directly found in the text Sun et al. (2022). Recently, EAE has been formulated as a generative task where pre-trained language models are guided with natural language to fill templates or generate arguments Hsu et al. (2022). Sharif et al. (2024) argue that this generative formulation better suits real-world applications as it can capture implicit and scattered arguments better. With the emergence of LLMs, generative model-based argument extraction gained more traction Sun et al. (2024); He et al. (2024); Gatto et al. (2025). So, we focus on generative extraction covering diverse models and datasets.
Evaluations for Generative EAE: Existing works for generative EAE primarily rely on exact matching for evaluation Huang et al. (2024). This strict approach unfairly penalizes models, even when the generated output is correct Fane et al. (2025). To address this, Han et al. (2024) adopts a relaxed matching approach, considering arguments similar if their embedding-based similarly exceeds a threshold of 0.5. Similarly, Sharif et al. (2024) used a threshold of 0.75. However, this approach has limitations. It fails to capture semantically similar arguments with different lexical or syntactic forms and wrongly classifies arguments with high token overlap as similar. Thus, performance reported solely on relaxed matching is unreliable. More recently, Lu et al. (2024) employed LLMs to determine argument similarity. Nonetheless, this approach incurs significant computational overhead and demands extensive human validation. Our REGen framework combines the strengths of exact, relaxed, and LLM-based matching. It systematically reduces misjudgments, computational costs, and the need for human validation.
@@ 7 Conclusion
This paper presents REGen, a novel evaluation framework for EAE. Our extensive experiments and human validation demonstrate its effectiveness, with a 23.93-point gain in average F1 score across six EAE datasets, and reliability, with 87.67% alignment with human judgments. We highlight the limitations of current evaluation approaches and illustrate how REGen addresses these issues. Furthermore, our analysis reveals that previous studies have underestimated the true performance of LLMs. We believe that REGen fills a critical gap in EAE research and motivates future work to explore the generative model’s capability in solving other information extraction tasks, e.g., relation extraction, entity extraction, and beyond.
@@ 8 Limitations
Intended Use: We released our judge and alignment datasets to facilitate future research on generative argument extraction evaluation.
Annotation: Judge and alignment data annotation were conducted by trained NLP researchers. All annotators were compensated as per the standard paying rate of the author’s institution. Key characteristics of our annotators include: (a) graduate students, (b) 3-6 years of research experience, and (c) a mix of native and non-native English speakers. We provided annotators detailed annotation guidelines, including argument extraction, semantical similarity, and the type of information we wanted to compare to mitigate potential biases.
Reproducibility Details on models and dataset processing are provided in Appendices C and D. The evaluation framework, code, and processed datasets are available at https://github.com/Omar-Sharif/REGen..
Acknowledgments
This work was supported in part by National Institutes of Health (NIH) grant 1R21DA059665-01A1. We also thank the reviewers for their valuable feedback.